OpenStudy (anonymous):

PLEASE HELP! Standard form to vertex form: y=-x^2-12x-35

OpenStudy (sidsiddhartha):

try to make a square first

hero (hero):

The vertex form of quadratic function has the form: y = a(x - h)^2 + k Where (h , k) is the vertex

hero (hero):

I can walk you through the steps to convert it if you like.

OpenStudy (anonymous):

Yes please! @Hero

hero (hero):

First, I would factor out the negative: y = -(x^2 + 12x + 35) Then divide both sides by -1 to get -y = x^2 + 12x + 35 Next I observe that (b/2)^2 = (12/2)^2 = (6)^2 = 36 The 36 is needed in order to be able to write the quadratic in square form. Notice that 35 = 36 - 1 so -y = x^2 + 12x + 36 - 1 Now we can write x^2 + 12x + 36 as (x + 6)^2 since it is in square form: -y = (x + 6)^2 - 1 Now divide the both sides by -1 again to place the equation in proper form: y = -(x + 6)^2 + 1

OpenStudy (anonymous):

Start by 'completing the square' Step 1) take the coefficient if the 'x' term and divide it by 2 Step 2) square the value u got from step 1 Step 3) insert that value between the 'x' term and the constant in your equation and subtract the same value from the constant on the end. Good so far?

OpenStudy (anonymous):

Thank you so very much! @Hero